# SafeSieve: From Heuristics to Experience in Progressive Pruning for Multi-Agent LLM Communication

### v1

@@ Abstract
LLM-based multi-agent systems exhibit strong collaborative capabilities but often suffer from redundant communication and excessive token overhead. Existing methods typically enhance efficiency through pretrained GNNs or greedy algorithms, but often isolate pre- and post-task optimization, lacking a unified strategy. To this end, we present SafeSieve, a progressive and adaptive multi-agent pruning algorithm that dynamically refines the inter-agent communication through a novel dual-mechanism. SafeSieve integrates initial LLM-based semantic evaluation with accumulated performance feedback, enabling a smooth transition from heuristic initialization to experience-driven refinement. Unlike existing greedy Top-k pruning methods, SafeSieve employs 0-extension clustering to preserve structurally coherent agent groups while eliminating ineffective links. Experiments across benchmarks (SVAMP, HumanEval, etc.) showcase that SafeSieve achieves 94.01% average accuracy while reducing token usage by 12.4%-27.8%. Results further demonstrate robustness under prompt injection attacks (1.23% average accuracy drop). In heterogeneous settings, SafeSieve reduces deployment costs by 13.3% while maintaining performance. These results establish SafeSieve as a robust, efficient, and scalable framework for practical multi-agent systems. Our code can be found in https://anonymous.4open.science/r/SafeSieve-D8F2FFUN.
@@ Introduction
Large language model (LLM) based multi-agent systems (MAS) have demonstrated impressive collaborative problem-solving capabilities (Wang et al. 2025a; Chang et al. 2024), fueling frameworks such as AutoGen and ChatDev for real-world applications (Wu et al. 2023; Chen et al. 2023). Nevertheless, the dense, round-robin conversations among agents often incur substantial token overhead and communication redundancy, which not only elevates inference cost but also dilutes attention over key information, leading to potential accuracy degradation (Liu et al. 2024a). Longer context windows further enlarge the attack surface for prompt-injection (Anil et al. 2024). Consequently, recent studies have begun to sparsify MAS communication topologies to improve both efficiency and robustness (Zhuge et al. 2024; Zhang, Yue et al. 2024b, a).
@@
Among communication sparsification strategies, one prevalent line of work directly constructs compact graph topologies prior to execution, such as G-Designer and GPTSwarm (Zhang, Yue et al. 2024b; Zhuge et al. 2024). These methods improve communication efficiency at initialization but exhibit limited generalizability and cannot adapt to run-time dynamics. A more recent and adaptive paradigm is on-the-fly pruning, where the communication graph is dynamically adjusted based on task performance feedback. Representative works like AgentPrune and AgentDropout (Zhang, Yue et al. 2024a; Wang, Wang et al. 2025) start from a basic or fully-connected topology and iteratively prune edges during execution. These approaches require no pretraining and offer strong task adaptability with minimal deployment burden. However, most of them still rely on greedy top-$k$ pruning strategies, which may mistakenly remove critical communication paths, reducing system robustness. To date, no method has yet unified both heuristic-driven early filtering and performance-aware dynamic adaptation, forming a full-spectrum optimization pipeline.
Motivated by the similarity between MAS collaboration and human team organization (Guo et al. 2024), we propose SafeSieve, a progressive and adaptive pruning algorithm. SafeSieve introduces a two-stage edge scoring scheme that ➀ uses LLM-based semantic compatibility to offer heuristic guidance at startup and ➁ gradually shifts weight to accumulated contribution during execution, emulating the “plan-then-adjust” paradigm of human teamwork, as shown in Figure 1. Instead of pruning edges individually, SafeSieve employs a 0-extension based clustering mechanism (Fakcharoenphol et al. 2003), preserving structurally coherent agent groups while eliminating ineffective links. This design avoids the local sub-optimality of greedy top-$k$ pruning and retains inter-agent complementarity.
@@
Pre-design methods, including GPTSwarm (Zhuge et al. 2024), G-Designer (Zhang, Yue et al. 2024b), AnyMAC (Wang et al. 2025b), EvoMAC (Hu et al. 2024), and DyLAN (Liu et al. 2024c), focus on directly generating efficient communication topologies—whether through GNN-based graph construction, autoregressive agent selection, evolutionary adaptation, or two-stage team formation.
Post-pruning approaches begin with dense communication topology and progressively sparsify them: AgentPrune (Zhang, Yue et al. 2024a) introduces dynamic edge pruning via one-hot mask matrices, AgentDropout (Wang, Wang et al. 2025) extends this to node-level pruning, Adaptive Graph Pruning (Li et al. 2025) jointly learns node selection and edge connectivity via end-to-end GNN training, and Adaptive Prompt Pruning (Dong, Chen, and Chi 2024) reduces per-agent prompt lengths to save token.
@@
Recent studies have shown that LLMs exhibit human-like collaborative patterns, where agents with similar or complementary capabilities naturally form effective working groups (Wang et al. 2024; Hong et al. 2023). Inspired by this observation, we propose that clustering-based pruning offers a more principled approach than direct edge removal in MAS collaboration.
While various clustering methods exist—including spectral clustering (Schaeffer 2007), hierarchical clustering (Xue et al. 2024), and density-based approaches (Birant and Kut 2007)—the $k$-terminal 0-extension problem (Calinescu, Karloff, and Rabani 2003; Fakcharoenphol et al. 2003) presents unique advantages for our setting. 0-extension has been successfully applied in computing system for its computational efficiency ($O(n\log n)$ complexity), simple deployment, and strong theoretical guarantees in preserving graph connectivity (Englert et al. 2014; Chen and Gopalakrishnan 1998). By formulating agent clustering as a 0-extension problem, we replace aggressive top-$k$ pruning with a connectivity-aware approach that maintains critical communication paths between agent communities while achieving similar sparsification rates.
@@
Building upon the graph-based paradigm for multi-agent systems, GPTSwarm (Zhuge et al. 2024) first formalized multi-agent orchestration as a differentiable computational graph. At each communication round $t$, the interaction among $N$ agents is represented as a directed communication graph $\mathcal{G}_{t}=(\mathcal{V},\mathcal{E}_{t})$, where $\mathcal{V}$ denotes the set of agents and each edge $e_{ij}\in\mathcal{E}_{t}$ represents a message from agent $i$ to agent $j$. While GPTSwarm generates static graph structures during inference, AgentPrune (Zhang, Yue et al. 2024a) introduces dynamic sparsification through a mask matrix $\mathbf{M}^{(t)}\in\{0,1\}^{N\times N}$, where each entry controls edge activation:
@@
This mask effectively defines a candidate set of communication links from which the actual runtime communication graph is sampled. AgentDropout (Wang, Wang et al. 2025) extends this framework by incorporating node-level pruning and real-time feedback mechanisms, enabling more aggressive sparsification across communication rounds.
@@
While semantic similarity facilitates basic cooperation (Deng et al. 2025), functional complementarity plays a more critical role in complex multi-hop reasoning tasks (Zhang et al. 2024; Zhang, Chen, and Kumar 2025). SafeSieve initializes the communication edge score between agents $i$ and $j$ by combining embedding-based similarity with expert-assessed compatibility:
@@
To capture the dynamic nature of agent cooperation, SafeSieve progressively shifts from static semantic initialization to experience-based scoring through a unified temporal mechanism (Zhang, Yang, and Başar 2022).
@@
SafeSieve employs a principled clustering approach based on the 0-extension framework (Fakcharoenphol et al. 2003) to make globally-informed pruning decisions. This method provides theoretical approximation guarantees while maintaining computational and communicative efficiency.
@@
$\theta(t)=\theta_{0}+(\theta_{\max}-\theta_{0})\cdot\left[1-\exp^{-k\cdot\max\left(\frac{t}{T}-1,0\right)}\right]$ (5)
@@
Table 1: Performance comparison between SafeSieve and baseline reasoning frameworks. Results for Vanilla and CoT under DeepSeek-V3 are adapted from AgentDropout (Wang, Wang et al. 2025) except MMLU and MATH-500; GPT-4o-mini results are taken from AGP (Li et al. 2025) except for MATH-500 and post-prune paradigm. Other results are evaluated by us under the same computing environment. All methods that start with a basic topology are based on full-connected graph.
@@
In our main experiments, we adopt Deepseek-V3 (671B) (Liu et al. 2024b) as the primary backbone model. For smaller-model ablation, we use GPT-4o-mini. In heterogeneous settings, we additionally incorporate LLaMA3-8B (AI@Meta 2024), Qwen2.5-72B (Team 2024), and Kimi-K2 (AI 2024) to validate cross-model communication robustness. Build upon these, we evaluate general and mathematical reasoning using six standard benchmarks: MMLU (Hendrycks et al. 2021), GSM8K (Cobbe et al. 2021), SVAMP (Patel, Bhattamishra, and Goyal 2021), HumanEval (Chen et al. 2021), AQuA (Ling et al. 2017) and MATH-500 (Lightman et al. 2023).
@@
In the main experiments, we compare SafeSieve with prompting-based strategies including Vanilla (direct reasoning) and Chain-of-Thought (CoT) (Wei et al. 2023)(referred to as single), as well as collaborative frameworks such as GPT-Swarm (Zhuge et al. 2024) and G-Designer (Zhang, Yue et al. 2024b), which enhance agent outputs prior to inference (referred to as pre-design). We also include pruning-based baselines AgentPrune (Zhang, Yue et al. 2024a) and AgentDropout (Wang, Wang et al. 2025), which dynamically remove redundant links in multi-agent communication graphs (referred to as post-prune). Since SafeSieve also belongs to the pruning paradigm, these two methods are further compared under prompt injection and heterogeneous-agent settings.
@@
Takeaway ➊: Carefully designed communication graphs outperform single-agent methods, with post-pruning paradigms comprehensively surpassing pre-design approaches.
@@
Takeaway ➋: Task-dependent performance improvements exhibit differentiated characteristics.
@@
Takeaway ➌: Large and small models show differentiated improvements across task types, providing necessity for heterogeneous deployment.
@@
Takeaway ➍: SafeSieve achieves superior efficiency-accuracy trade-off positioning across all benchmarks.
@@
Takeaway ➎: Ablation experiments validate the effectiveness of three core components.
@@
Takeaway ➊: Task characteristics determine vulnerability patterns to malicious agents.
@@
Takeaway ➋: SafeSieve’s triple defense mechanism ensures minimal degradation and superior robustness.
@@
Takeaway ➊: Large model commander effect significantly reduces system costs.
@@
Takeaway ➋: Task-dependent heterogeneous effects exhibit barrel principle characteristics.
@@ Conclusion
We propose SafeSieve, a principled pruning framework for multi-agent LLM collaboration that unifies semantic initialization with experience-guided refinement. By jointly leveraging prompt-based similarity and outcome-driven complementarity, SafeSieve dynamically sparsifies the communication topology through 0-extension clustering. Experimental results across six benchmarks—including reasoning, coding, and knowledge tasks—demonstrate that SafeSieve consistently outperforms both prompt-only and existing multi-agent collaboration baselines, achieving up to 6.58% accuracy improvement and 30% token reduction. Moreover, it exhibits strong robustness under agent injection and excels in heterogeneous settings where models vary in scale and capability. These results validate the effectiveness of structure-aware, behavior-adaptive pruning in enabling scalable, efficient, and reliable LLM cooperation.

### v2

# SafeSieve: From Heuristics to Experience in Progressive Pruning for LLM-based Multi-Agent Communication

@@ Abstract
LLM-based multi-agent systems exhibit strong collaborative capabilities but often suffer from redundant communication and excessive token overhead. Existing methods typically enhance efficiency through pretrained GNNs or greedy algorithms, but often isolate pre- and post-task optimization, lacking a unified strategy. To this end, we present SafeSieve, a progressive and adaptive multi-agent pruning algorithm that dynamically refines the inter-agent communication through a novel dual-mechanism. SafeSieve integrates initial LLM-based semantic evaluation with accumulated performance feedback, enabling a smooth transition from heuristic initialization to experience-driven refinement. Unlike existing greedy Top-k pruning methods, SafeSieve employs 0-extension clustering to preserve structurally coherent agent groups while eliminating ineffective links. Experiments across benchmarks (SVAMP, HumanEval, etc.) showcase that SafeSieve achieves 94.01% average accuracy while reducing token usage by 12.4%-27.8%. Results further demonstrate robustness under prompt injection attacks (1.23% average accuracy drop). In heterogeneous settings, SafeSieve reduces deployment costs by 13.3% while maintaining performance. These results establish SafeSieve as an efficient, GPU-free, and scalable framework for practical multi-agent systems. Our code can be found below.
Codehttps://github.com/csgen/SafeSieve
@@ Introduction
Large language model (LLM) based multi-agent systems (MAS) have demonstrated impressive collaborative problem-solving capabilities (wang2025comprehensive; chang2024survey), fueling frameworks such as AutoGen and ChatDev for real-world applications (AutoGen2023; chatdev). Nevertheless, the dense, round-robin conversations among agents often incur substantial token overhead and communication redundancy, which not only elevates inference cost but also dilutes attention over key information, leading to potential accuracy degradation (Liu2023LostMid). Longer context windows further enlarge the attack surface for prompt-injection (anil2024manyshot). Consequently, recent studies have begun to sparsify MAS communication topologies to improve both efficiency and robustness (zhuge2024gptswarm; Zhang2024GDesigner; Zhang2024AgentPrune).
@@
Among communication sparsification strategies, one prevalent line of work directly constructs compact graph topologies prior to execution, such as G-Designer and GPTSwarm (Zhang2024GDesigner; zhuge2024gptswarm). These methods improve communication efficiency at initialization but exhibit limited generalizability and cannot adapt to run-time dynamics. A more recent and adaptive paradigm is on-the-fly pruning, where the communication graph is dynamically adjusted based on task performance feedback. Representative works like AgentPrune and AgentDropout (Zhang2024AgentPrune; Wang2025AgentDropout) start from a basic or fully-connected topology and iteratively prune edges during execution. These approaches require no pretraining and offer strong task adaptability with minimal deployment burden. However, most of them still rely on greedy top-$k$ pruning strategies, which may mistakenly remove critical communication paths, reducing system robustness. To date, no method has yet unified both heuristic-driven early filtering and performance-aware dynamic adaptation, forming a full-spectrum optimization pipeline.
Motivated by the similarity between MAS collaboration and human team organization (guo2024embodied_llm_teams), we propose SafeSieve, a progressive and adaptive pruning algorithm. SafeSieve introduces a two-stage edge scoring scheme that ➀ uses LLM-based semantic compatibility to offer heuristic guidance at startup and ➁ gradually shifts weight to accumulated contribution during execution, emulating the “plan-then-adjust” paradigm of human teamwork, as shown in Figure 1. Instead of pruning edges individually, SafeSieve employs a 0-extension based clustering mechanism (fakcharoenphol2003improved0extension), preserving structurally coherent agent groups while eliminating ineffective links. This design avoids the local sub-optimality of greedy top-$k$ pruning and retains inter-agent complementarity.
@@
Pre-design methods, including GPTSwarm (zhuge2024gptswarm), G-Designer (Zhang2024GDesigner), AnyMAC (wang2025anymac), EvoMAC (hu2024selfevolving), and DyLAN (liu2024dynamicllm), focus on directly generating efficient communication topologies—whether through GNN-based graph construction, autoregressive agent selection, evolutionary adaptation, or two-stage team formation.
Post-pruning approaches begin with dense communication topology and progressively sparsify them: AgentPrune (Zhang2024AgentPrune) introduces dynamic edge pruning via one-hot mask matrices, AgentDropout (Wang2025AgentDropout) extends this to node-level pruning, Adaptive Graph Pruning (li2025adaptivegraph) jointly learns node selection and edge connectivity via end-to-end GNN training, and Adaptive Prompt Pruning (dong2024promptprompted) reduces per-agent prompt lengths to save token.
@@
Recent studies have shown that LLMs exhibit human-like collaborative patterns, where agents with similar or complementary capabilities naturally form effective working groups (wang2023unleashing; hong2023metagpt). Inspired by this observation, we propose that clustering-based pruning offers a more principled approach than direct edge removal in MAS collaboration.
While various clustering methods exist—including spectral clustering (schaeffer2007graph), hierarchical clustering (xue2024comprehensive), and density-based approaches (birant2007stdbscan)—the $k$-terminal 0-extension problem (calinescu2003approximation; fakcharoenphol2003improved0extension) presents unique advantages for our setting. 0-extension has been successfully applied in computing system for its computational efficiency ($O(n\log n)$ complexity), simple deployment, and strong theoretical guarantees in preserving graph connectivity (englert2014vertex; chen2011clustering). By formulating agent clustering as a 0-extension problem, we replace aggressive top-$k$ pruning with a connectivity-aware approach that maintains critical communication paths between agent communities while achieving similar sparsification rates.
@@
Building upon the graph-based paradigm for multi-agent systems, GPTSwarm (zhuge2024gptswarm) first formalized multi-agent orchestration as a differentiable computational graph. At each communication round $t$, the interaction among $N$ agents is represented as a directed communication graph $\mathcal{G}_{t}=(\mathcal{V},\mathcal{E}_{t})$, where $\mathcal{V}$ denotes the set of agents and each edge $e_{ij}\in\mathcal{E}_{t}$ represents a message from agent $i$ to agent $j$. While GPTSwarm generates static graph structures during inference, AgentPrune (Zhang2024AgentPrune) introduces dynamic sparsification through a mask matrix $\mathbf{M}^{(t)}\in\{0,1\}^{N\times N}$, where each entry controls edge activation:
@@
This mask effectively defines a candidate set of communication links from which the actual runtime communication graph is sampled. AgentDropout (Wang2025AgentDropout) extends this framework by incorporating node-level pruning and real-time feedback mechanisms, enabling more aggressive sparsification across communication rounds.
@@
While semantic similarity facilitates basic cooperation (deng2025semantic_multi_agent_comm), functional complementarity plays a more critical role in complex multi-hop reasoning tasks (zhang2024chain_of_agents; zhang2025multi_agent_collaboration_survey). SafeSieve initializes the communication edge score between agents $i$ and $j$ by combining embedding-based similarity with expert-assessed compatibility:
@@
To capture the dynamic nature of agent cooperation, SafeSieve progressively shifts from static semantic initialization to experience-based scoring through a unified temporal mechanism (zhang2021multiagent_deep_rl).
@@
SafeSieve employs a principled clustering approach based on the 0-extension framework (fakcharoenphol2003improved0extension) to make globally-informed pruning decisions. This method provides theoretical approximation guarantees while maintaining computational and communicative efficiency.
@@
$\theta(t)=\theta_{0}+(\theta_{\max}-\theta_{0})\cdot\left[1-\exp^{-k\cdot\max\left(\frac{t}{T},0\right)}\right]$ (5)
@@
Table 1: Performance comparison between SafeSieve and baseline reasoning frameworks. Results for Vanilla and CoT under DeepSeek-V3 are adapted from AgentDropout (Wang2025AgentDropout) except MMLU and MATH-500; GPT-4o-mini results are taken from AGP (li2025adaptivegraph) except for MATH-500 and post-prune paradigm. Other results are evaluated by us under the same computing environment. All methods that start with a basic topology are based on full-connected graph.
@@
In our main experiments, we adopt Deepseek-V3 (671B) (Liu2024Deepseek) as the primary backbone model. For smaller-model ablation, we use GPT-4o-mini. In heterogeneous settings, we additionally incorporate LLaMA3-8B (llama3modelcard), Qwen2.5-72B (qwen2.5), and Kimi-K2 (Moonshot2024Kimi) to validate cross-model communication robustness. Build upon these, we evaluate general and mathematical reasoning using six standard benchmarks: MMLU (hendryckstest2021), GSM8K (cobbe2021gsm8k), SVAMP (patel-etal-2021-nlp), HumanEval (chen2021evaluating), AQuA (ling2017program) and MATH-500 (lightman2023lets).
@@
In the main experiments, we compare SafeSieve with prompting-based strategies including Vanilla (direct reasoning) and Chain-of-Thought (CoT) (wei2023chainofthoughtpromptingelicitsreasoning)(referred to as single), as well as collaborative frameworks such as GPT-Swarm (zhuge2024gptswarm) and G-Designer (Zhang2024GDesigner), which enhance agent outputs prior to inference (referred to as pre-design). We also include pruning-based baselines AgentPrune (Zhang2024AgentPrune) and AgentDropout (Wang2025AgentDropout), which dynamically remove redundant links in multi-agent communication graphs (referred to as post-prune). Since SafeSieve also belongs to the pruning paradigm, these two methods are further compared under prompt injection and heterogeneous-agent settings.
@@
Takeaway 1: Carefully designed communication graphs outperform single-agent methods, with post-pruning paradigms comprehensively surpassing pre-design approaches.
@@
Takeaway 2: Task-dependent performance improvements exhibit differentiated characteristics.
@@
Takeaway 3: Large and small models show differentiated improvements across task types, providing necessity for heterogeneous deployment.
@@
Takeaway 4: SafeSieve achieves superior efficiency-accuracy trade-off positioning across all benchmarks.
@@
Takeaway 5: Ablation experiments validate the effectiveness of three core components.
@@
Takeaway 1: Task characteristics determine vulnerability patterns to malicious agents.
@@
Takeaway 2: SafeSieve’s triple defense mechanism ensures minimal degradation and superior robustness.
@@
Takeaway 1: Large model commander effect significantly reduces system costs.
@@
Takeaway 2: Task-dependent heterogeneous effects exhibit barrel principle characteristics.
@@ Conclusion
We propose SafeSieve, a principled pruning framework for multi-agent collaboration that unifies semantic initialization with experience-guided refinement. It provides GPU-free sparsing strategy. Experiments across six benchmarks, including reasoning and coding tasks, show SafeSieve outperforms baselines with up to 6.58% accuracy gains and 30% token reduction. Furthermore, it demonstrates robustness against agent injection and excels in heterogeneous settings varying in model scale, validating the efficacy of structure-aware pruning for efficient LLM cooperation.